Fermi National Laboratory


Accelerator Update

Monday March 29
The day shift began with Operations monitoring store 3328, with the Antiproton Source (Pbar) stacking, with experts tuning the AP1 transfer line and Booster, and with MiniBooNE and Switchyard (SY) taking beam.

A Booster expert began a study period at 9:40 AM.

At 11:39 AM, the Main Injector (MI) RF area vacuum valves closed due to vacuum pumps tripping off. Operators got everything back on and opened.

Operations reported at 12:28 PM that the SY Septa for Meson beam had been tripping off during the morning. It continued to do this all afternoon. A SY expert said there was a vacuum problem and it would be repaired within the next few days. Since only MCenter was taking beam, the Septa could be left off.

A MI expert began a 27 GeV slow acceleration study period at 6:50 PM.

At 7:55 PM, operators noticed that a Tevatron (TeV) quadrupole power supply (DQ9) had tripped off due to a quench indication, which aborted store 3328. Operators intentionally turned off the TeV to avoid a larger quench at sector D1. Actually, the magnet had not quenched. Operators contacted power supply technicians.

Techs managed to get DQ9 back on at 9:26 PM.

Operations had put the TeV into a dry squeeze, but at 10:21 PM another quadrupole power supply (B0Q3) tripped off due to a blown fuse. Operations contacted an expert.

Tuesday March 30
The midnight shift began with Pbar stacking, with MiniBooNE and SY taking beam, and with techs investigating B0Q3 problems.

A TeV RF station (TRF1) tripped off at 3:33 AM. Operators investigated and discovered a bad fan, which they unplugged and reset the station.

Techs reported at 4 AM that a bad fan had caused a problem in the 110-volt circuit for B0Q3.

A TeV expert asked Operations at 4:31 AM to put the TeV into an hour-long dry squeeze to make sure the low beta magnets stay on.

At 8:15 AM, a Pbar Accumulator RF station (ARF1) tripped off. It reset okay.

Operations put the TeV into shot setup at 8:41 AM. At 11:53 AM, Operations established store 3335 with an initial luminosity of 50.47E30.

At 12:02 PM, experts worked at tuning the TeV beam. The Collider Detector at Fermilab (CDF) had to turn off some of its detectors due to this problem. Experts discovered at 1:45 PM that a vertical collimator was misaligned. Realigning this collimator helped solve the tuning problem and CDF was able to turn back on.

Pbar resumed stacking at 12:38 PM.

Operations reported at 4:15 PM that for the past hour MI had been having serious beam problems. A MI expert said it was a Booster problem. A Booster expert traced the problem to an injection phase offset from Booster. MI beam was back to normal by 5:22 PM.

Operators began sending beam to MCenter at 6:15 PM. Experts tuned the beam.

A TeV RF station (TRF8) tripped off at 11:51 PM on an overcurrent indication. It was left off.

Wednesday March 31
The midnight shift began with Operations monitoring store 3335, with Pbar stacking, and with MiniBooNE and SY/MCenter taking beam.

Operations turned off the SY/MCenter beam at 7:06 AM for Septa vacuum work.

Today's Plans
The plans for today includes a two to three hour access this morning into MI and Booster and then to establish a new store this afternoon.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 3/31/2004   email Fermilab